snowycatt1527:

For dinner, Abelita chooses the Italian salad that costs $3.99 before tax, and Franco chooses the fettuccini alfredo that costs $7.17 before tax. If the sales tax is 5 percent, explain how to find the total cost for their dinner.

jimthompson5910:

Here's a similar problem: A person orders salad for $5 and Fettuccini Alfredo for $7. Both of these costs are before tax. If sales tax is 8%, then what is the total bill? Ignore any tips.


Solution: Salad = $5 before tax Fettuccini Alfredo = $7 before tax total cost before tax = 5+7 = 12 dollars A sales tax of 8% means 8% of 12 = 0.08*12 = 0.96 dollars is the tax The total bill comes to 12+0.96 = 12.96 This is without considering any tip A shortcut is to say 12*1.08 = 12.96 The 1.08 multiplier means "increase by 8%" Keep in mind that this is an example and the answer 12.96 doesn't apply to your question exactly, but you'll use these same steps to get the answer you're after.